OpenStudy (anonymous):

what functions are second derivative of itself? what functions are negative second derivative?

OpenStudy (ghazi):

exponential function are second derivative of itself

OpenStudy (ghazi):

\[\frac{ d(e^x) }{ dt }=e^x\]

OpenStudy (cwrw238):

yup

OpenStudy (eyust707):

trig functions or negative second derivatives of themselves

OpenStudy (eyust707):

\[{d^2 \over dx^2 } \sin x = - sinx \]